OpenStudy (anonymous):

Evaluate the indefinite integral: sqrt[10x-(x^2)]dx

zepdrix (zepdrix):

Hey c:\[\Large\bf\sf \int\limits \sqrt{10x-x^2}\;dx\]

zepdrix (zepdrix):

So ummmmmmm, We need to complete the square and all that business I guess.

zepdrix (zepdrix):

Might be easier to work with if you factor the negative out of each term first.\[\Large\bf\sf \int\limits\limits \sqrt{-(x^2-10x)}\;dx\]So we want to complete the square on the stuff in the brackets. Remember how to do that? :O Half of the b coefficient and square it.